# ROME: Robust Multi-Modal Density Estimator [^1] Extended version with Appendix: https://arxiv.org/abs/2401.10566

### v2

@@ Abstract
The estimation of probability density functions is a fundamental problem in science and engineering. However, common methods such as kernel density estimation (KDE) have been demonstrated to lack robustness, while more complex methods have not been evaluated in multi-modal estimation problems. In this paper, we present ROME (RObust Multi-modal Estimator) [^3] Source code: https://github.com/anna-meszaros/ROME, a non-parametric approach for density estimation which addresses the challenge of estimating multi-modal, non-normal, and highly correlated distributions. ROME utilizes clustering to segment a multi-modal set of samples into multiple uni-modal ones and then combines simple KDE estimates obtained for individual clusters in a single multi-modal estimate. We compared our approach to state-of-the-art methods for density estimation as well as ablations of ROME, showing that it not only outperforms established methods but is also more robust to a variety of distributions. Our results demonstrate that ROME can overcome the issues of over-fitting and over-smoothing exhibited by other estimators.
@@ 1 Introduction
Figure 1: a) A schematic of a probabilistic model evaluation pipeline, using different metrics. b) ROME takes predicted samples of a given model and estimates their densities to enable the calculation of probability-based metrics (e.g., negative log-likelihood, Jensen-Shannon divergence, etc.).
@@ 3 RObust Multi-modal Estimator (ROME)
A solution to the above problem would be an estimator $f:\mathbb{R}^{N\times M}\rightarrow\mathcal{P}$, resulting in $\widehat{p}=f(\bm{X})$. Our proposed estimator $f_{\text{ROME}}$ (Algorithm 1) is built on top of non-parametric cluster extraction. Namely, by separating groups of samples surrounded by areas of low density – expressing the mode of the underlying distribution – we reduce the multi-modal density estimation problem to multiple uni-modal density estimation problems for each cluster. The distributions within each cluster then become less varied in density or correlation than the full distribution. Combining this with decorrelation and normalization, the use of established methods such as KDE to estimate probability densities for those uni-modal distributions is now more promising, as problems with multi-modality and correlated modes (see Section 2) are accounted for. The multi-modal distribution is then obtained as a weighted average of the estimated uni-modal distributions.
@@ 3.1 Extracting Clusters
where ${\min}_{k_{\bm{c}}}$ is the $k_{\bm{c}}$-smallest value from all the available $\bm{\widehat{x}}$, which is used to smooth out random local density fluctuations. We use
@@
where $k_{\min}$ and $k_{\max}$ ensure that there are always sufficient but not too many points used for this smoothing. Meanwhile, the term $NM/\alpha_{k}$ adjusts $k_{\bm{c}}$ to the number of samples and dimensions.
@@ 3.3 Normalization
Figure 2: Samples from the two-dimensional synthetic distributions used for evaluating different probability density estimators.
@@ 3.4 Estimating the Probability Density Function
Distribution ROME MPW VC ROME MPW VC ROME MPW VC
@@ 5.1 Baseline Comparison
Norm. No norm. $f_{\text{GMM}}$
Clustering Decorr. No decorr.
Silhouette $13\scriptscriptstyle{\pm 0.20}- 0$ $13\scriptscriptstyle{\pm 0.20}- 0$ $13\scriptscriptstyle{\pm 0.19}- 0$ $14\scriptscriptstyle{\pm 0.20}- 0$
[…]
@@
Clustering Decorr. No decorr.
@@
Clustering Decorr. No decorr.
@@
No clusters $1.82\scriptscriptstyle{\pm 0.60}$ $1.84\scriptscriptstyle{\pm 0.60}$ $3.17\scriptscriptstyle{\pm 0.89}$ $5.29\scriptscriptstyle{\pm 1.34}$
@@
Figure 5: Samples generated by ROME and KDE – equivalent to ROME without clustering, decorrelation and normalization – (in pink) contrasted with samples from $p$ (in blue); Aniso. Note that the samples by KDE are more spread out which indicates over-smoothing.
@@ 5.2 Ablation Studies
Testing variants of ROME on the Aniso distribution (Table 3) demonstrated not only the need for decorrelation, through the use of rotation, but also normalization in the case of distributions with highly correlated features. There, using either of the two clustering methods in combination with normalization and decorrelation (our full proposed method) is better than the two alternatives of omitting only decorrelation or both decorrelation and normalization. In the case of clustering with the silhouette score, the full method is significantly more likely to reproduce the underlying distribution $p$ by a factor of $1.19$ (with a statistical significance of $10^{-50}$, see Appendix B) as opposed to omitting only decorrelation, and by $1.30$ compared to omitting both decorrelation and normalization. Similar trends can be seen when clustering based on DBCV, with the full method being more likely to reproduce $p$ by a factor of $1.14$ and $1.31$ respectively. Results on the Aniso distribution further show that KDE on its own is not able to achieve the same results as ROME, but rather it has a tendency to over-smooth (Figure 5). Additionally, the ablation with and without normalization on the Two Moons distribution (Table 4) showed that normalization is necessary to avoid over-smoothing in the case of non-normal distributions.
@@
Figure 6: Samples generated by ROME, and ROME with $f_{\text{GMM}}$ as the downstream estimator (in pink) contrasted with samples from $p$ (in blue); Two Moons. Note that the samples when using $f_{\text{GMM}}$ are more spread out, despite having a component for each of the two clusters, which clearly displays over-smoothing.
@@ 6 Conclusion
In our comparison against two established and sophisticated density estimators, we observed that ROME achieved consistently good performance across all test cases, while manifold Parzen windows (MPW) and vine copulas (VC) were susceptible to over-fitting and over-smoothing. For example, while MPW is superior at capturing non-normal distributions, it produces kernels with too small of a bandwidth (hence the over-fitting), which is likely caused by the selected number of nearest neighbors used for the localized kernel estimation being too small. Meanwhile, compared to VC, ROME is numerically more stable and does not hallucinate new modes in the estimated densities (Figure 4). Furthermore, as part of several ablation studies, we found that ROME overcomes the shortcomings of other common density estimators, such as the over-fitting exhibited by kNN or the over-smoothing by GMM. In those studies, we additionally demonstrated that our approach of using clustering, decorrelation, and normalization is indispensable for overcoming the deficiencies of KDE.
@@ C Clustering Performance
$f_{\text{ROME}}$ $f_{\text{MPW}}$ $f_{\text{VC}}$
Aniso $\underline{0.007}^{\pm 0.001}$ $0.016^{\pm 0.001}$ $0.143^{\pm 0.003}$
Varied $\underline{0.078}^{\pm 0.003}$ $0.084^{\pm 0.003}$ $0.087^{\pm 0.003}$
@@
Silhouette $0.007^{\pm 0.001}$ $0.014^{\pm 0.001}$ $0.039^{\pm 0.001}$ $0.020^{\pm 0.001}$ $0.054^{\pm 0.001}$ $0.020^{\pm 0.001}$ $0.001^{\pm 0.000}$
DBCV $0.009^{\pm 0.002}$ $0.014^{\pm 0.001}$ $0.037^{\pm 0.001}$ $0.022^{\pm 0.001}$ $0.061^{\pm 0.003}$ $0.023^{\pm 0.002}$ $0.008^{\pm 0.003}$
No clusters $0.095^{\pm 0.001}$ $0.015^{\pm 0.001}$ $0.245^{\pm 0.002}$ $0.022^{\pm 0.001}$ $0.039^{\pm 0.001}$ $0.023^{\pm 0.001}$ $0.530^{\pm 0.002}$
@@
Normalization No normalization $f_{\text{GMM}}$
Decorrelation No decorrelation
Silhouette $0.078^{\pm 0.003}$ $0.083^{\pm 0.003}$ $0.078^{\pm 0.003}$ $0.083^{\pm 0.003}$ $0.070^{\pm 0.003}$ $0.083^{\pm 0.003}$ $0.074^{\pm 0.003}$
DBCV $0.074^{\pm 0.003}$ $0.084^{\pm 0.003}$ $0.074^{\pm 0.003}$ $0.084^{\pm 0.003}$ $0.065^{\pm 0.003}$ $0.085^{\pm 0.003}$ $0.068^{\pm 0.003}$
No clusters $0.173^{\pm 0.004}$ $0.081^{\pm 0.003}$ $0.183^{\pm 0.004}$ $0.080^{\pm 0.003}$ $0.074^{\pm 0.003}$ $0.080^{\pm 0.003}$ $0.540^{\pm 0.003}$
@@
Gaussian $\underline{0.006}^{\pm 0.001}$ $0.023^{\pm 0.002}$ $0.007^{\pm 0.001}$ $-0.02^{\pm 0.09}$ $-0.26^{\pm 0.05}$ $\underline{-0.01}^{\pm 0.11}$ $\underline{-2.84}^{\pm 0.02}$ $-2.87^{\pm 0.02}$ $-2.85^{\pm 0.02}$
Elliptical $\underline{0.005}^{\pm 0.001}$ $0.023^{\pm 0.002}$ $0.007^{\pm 0.001}$ $0.05^{\pm 0.17}$ $-0.29^{\pm 0.08}$ $\underline{0.04}^{\pm 0.16}$ $\underline{-1.64}^{\pm 0.02}$ $-1.66^{\pm 0.02}$ $\underline{-1.64}^{\pm 0.02}$
Rot. Ellip. $0.005^{\pm 0.001}$ $0.023^{\pm 0.002}$ $\underline{0.003}^{\pm 0.001}$ $\underline{0.04}^{\pm 0.20}$ $-0.35^{\pm 0.11}$ $0.84^{\pm 0.32}$ $\underline{-1.42}^{\pm 0.02}$ $-1.45^{\pm 0.02}$ $-1.55^{\pm 0.01}$
Uni. Traj. $\underline{0.001}^{\pm 0.000}$ $0.003^{\pm 0.000}$ $0.801^{\pm 0.003}$ $1.69^{\pm 0.01}$ $2.33^{\pm 0.01}$ $\underline{1.60}^{\pm 0.11}$ $\underline{32.46}^{\pm 0.02}$ $28.20^{\pm 0.02}$ $-193.88^{\pm 17.5}$
@@
Gaussian $\underline{0.004}^{\pm 0.001}$ $0.012^{\pm 0.001}$ $\underline{0.004}^{\pm 0.001}$
Elliptical $\underline{0.004}^{\pm 0.001}$ $0.012^{\pm 0.001}$ $\underline{0.004}^{\pm 0.001}$
Rot. Elliptical $\underline{0.004}^{\pm 0.001}$ $0.013^{\pm 0.001}$ $0.029^{\pm 0.001}$
@@ E Comparison on Uni-modal Distribution
Silhouette $0.004^{\pm 0.001}$ $0.004^{\pm 0.001}$ $0.017^{\pm 0.001}$ $0.000^{\pm 0.000}$
DBCV $0.004^{\pm 0.001}$ $0.004^{\pm 0.001}$ $0.017^{\pm 0.001}$ $0.000^{\pm 0.000}$
No clusters $0.004^{\pm 0.001}$ $0.004^{\pm 0.001}$ $0.017^{\pm 0.001}$ $0.000^{\pm 0.000}$
@@
Silhouette $0.004^{\pm 0.001}$ $0.027^{\pm 0.001}$ $0.038^{\pm 0.001}$ $0.000^{\pm 0.000}$
DBCV $0.004^{\pm 0.001}$ $0.027^{\pm 0.001}$ $0.038^{\pm 0.001}$ $0.000^{\pm 0.000}$
No clusters $0.004^{\pm 0.001}$ $0.027^{\pm 0.001}$ $0.038^{\pm 0.001}$ $0.000^{\pm 0.000}$
@@
Silhouette $\hphantom{0}32.46^{\pm 0.02}$ $\hphantom{0}31.87^{\pm 0.10}$ $-13.85^{\pm 0.00}$ $\hphantom{0}28.61^{\pm 0.02}$
DBCV $\hphantom{0}32.46^{\pm 0.02}$ $\hphantom{0}31.87^{\pm 0.10}$ $-14.51^{\pm 1.96}$ $\hphantom{0}28.61^{\pm 0.02}$
No clusters $\hphantom{0}32.46^{\pm 0.02}$ $\hphantom{0}31.87^{\pm 0.10}$ $-13.85^{\pm 0.00}$ $\hphantom{0}28.61^{\pm 0.02}$
@@
Silhouette $0.013^{\pm 0.004}$ $0.166^{\pm 0.063}$ $0.000^{\pm 0.000}$ $0.005^{\pm 0.003}$
DBCV $0.013^{\pm 0.004}$ $0.166^{\pm 0.063}$ $0.005^{\pm 0.031}$ $0.005^{\pm 0.003}$
No clusters $0.013^{\pm 0.004}$ $0.166^{\pm 0.063}$ $0.000^{\pm 0.000}$ $0.005^{\pm 0.003}$
@@
Norm. No norm. $f_{\text{GMM}}$
Clustering Decorr. No decorr.
Silhouette $\hphantom{0}1.69^{\pm 0.01}$ $\hphantom{0}1.68^{\pm 0.02}$ $22.39^{\pm 0.10}$ $\hphantom{0}2.20^{\pm 0.03}$
DBCV $\hphantom{0}1.69^{\pm 0.01}$ $\hphantom{0}1.68^{\pm 0.02}$ $23.18^{\pm 2.38}$ $\hphantom{0}2.20^{\pm 0.03}$
No clusters $\hphantom{0}1.69^{\pm 0.01}$ $\hphantom{0}1.68^{\pm 0.02}$ $22.39^{\pm 0.09}$ $\hphantom{0}2.20^{\pm 0.03}$

### v3

@@ Abstract
The estimation of probability density functions is a fundamental problem in science and engineering. However, common methods such as kernel density estimation (KDE) have been demonstrated to lack robustness, while more complex methods have not been evaluated in multi-modal estimation problems. In this paper, we present ROME (RObust Multi-modal Estimator), a non-parametric approach for density estimation which addresses the challenge of estimating multi-modal, non-normal, and highly correlated distributions. ROME utilizes clustering to segment a multi-modal set of samples into multiple uni-modal ones and then combines simple KDE estimates obtained for individual clusters in a single multi-modal estimate. We compared our approach to state-of-the-art methods for density estimation as well as ablations of ROME, showing that it not only outperforms established methods but is also more robust to a variety of distributions. Our results demonstrate that ROME can overcome the issues of over-fitting and over-smoothing exhibited by other estimators.
@@ 1 Introduction
Figure 1: ROME takes samples from unknown distributions and estimates their densities to enable further downstream applications.
@@ 3 RObust Multi-modal Estimator (ROME)
A solution to the above problem would be an estimator $f:\mathbb{R}^{N\times M}\rightarrow\mathcal{P}$, resulting in $\widehat{p}=f(\bm{X})$. Our proposed estimator $f_{\text{ROME}}$ [^1] Source code: https://github.com/anna-meszaros/ROME (Algorithm 1) is built on top of non-parametric cluster extraction. Namely, by separating groups of samples surrounded by areas of low density – expressing the mode of the underlying distribution – we reduce the multi-modal density estimation problem to multiple uni-modal density estimation problems for each cluster. The distributions within each cluster then become less varied in density or correlation than the full distribution. Combining this with decorrelation and normalization, the use of established methods such as KDE to estimate probability densities for those uni-modal distributions is now more promising, as problems with multi-modality and correlated modes (see Section 2) are accounted for. The multi-modal distribution is then obtained as a weighted average of the estimated uni-modal distributions.
@@ 3.1 Extracting Clusters
where ${\min}_{k_{\bm{c}}}$ is the $k_{\bm{c}}$-smallest value of all the available $\bm{\widehat{x}}$, used to smooth out random local density fluctuations. We use
@@
where $k_{\min}$ and $k_{\max}$ ensure there are sufficient but not too many points for this smoothing, while the term $NM/\alpha_{k}$ adjusts $k_{\bm{c}}$ to the number of samples and dimensions.
@@ 3.3 Normalization
Figure 2: Samples from the two-dimensional synthetic distributions used for evaluating different PDF estimators.
@@ 3.4 Estimating the Probability Density Function
Distrib. ROME MPW VC ROME MPW VC ROME MPW VC
@@ 5.1 Baseline Comparison
Norm. No norm. $f_{\text{GMM}}$
Cluster. Decorr. No decorr.
Silhouette $-0.13\scriptscriptstyle{\pm 0.20}$ $-0.13\scriptscriptstyle{\pm 0.20}$ $-0.13\scriptscriptstyle{\pm 0.19}$ $-0.14\scriptscriptstyle{\pm 0.20}$
DBCV $-0.09\scriptscriptstyle{\pm 0.23}$ $-0.09\scriptscriptstyle{\pm 0.23}$ $-0.07\scriptscriptstyle{\pm 0.23}$ ${-0.03}\scriptscriptstyle{\pm 0.24}$
No clus. $\hphantom{-}2.28\scriptscriptstyle{\pm 0.72}$ $\hphantom{-}2.26\scriptscriptstyle{\pm 0.72}$ $-0.17\scriptscriptstyle{\pm 0.26}$   $\>\,10.53\scriptscriptstyle{\pm 2.54}$
@@
Cluster. Decorr. No decorr.
@@
Cluster. Decorr. No decorr.
@@
No clus. $1.82\scriptscriptstyle{\pm 0.60}$ $1.84\scriptscriptstyle{\pm 0.60}$ $3.17\scriptscriptstyle{\pm 0.89}$ $5.29\scriptscriptstyle{\pm 1.34}$
@@
Figure 5: Samples generated by ROME and KDE – equivalent to ROME without clustering, decorrelation and normalization – (in pink) contrasted with samples from $p$ (in blue); Aniso. Note that the samples by KDE are more spread out, indicating over-smoothing.
@@ 5.2 Ablation Studies
Testing variants of ROME on the Aniso distribution (Table 3) demonstrated not only the need for decorrelation, through the use of rotation, but also normalization in the case of distributions with highly correlated features. There, using either of the two clustering methods in combination with normalization and decorrelation (our full proposed method) is better than the two alternatives of omitting only decorrelation or both decorrelation and normalization. In the case of clustering with the silhouette score, the full method is significantly more likely to reproduce the underlying distribution $p$ by a factor of $1.19$ (with a statistical significance of $10^{-50}$, see Appendix B) as opposed to omitting only decorrelation, and by $1.30$ compared to omitting both decorrelation and normalization. Similar trends can be seen when clustering based on DBCV, with the full method being more likely to reproduce $p$ by a factor of $1.14$ and $1.31$ respectively. Results on the Aniso distribution further show that KDE on its own is not able to achieve the same results as ROME, but rather it has a tendency to over-smooth (Figure 5). Additionally, the ablation with and without normalization on the Two Moons distribution (Table 4) showed that normalization is necessary to avoid over-smoothing on non-normal distributions.
@@
Figure 6: Samples generated by ROME, and ROME with $f_{\text{GMM}}$ as the downstream estimator (in pink) contrasted with samples from $p$ (in blue); Two Moons. Note that the samples from $f_{\text{GMM}}$ are more spread out, which clearly displays over-smoothing.
@@ 6 Conclusion
In our comparison against two established and sophisticated density estimators, we observed that ROME achieved consistently good results across all tests, while manifold Parzen windows (MPW) and vine copulas (VC) were susceptible to over-fitting and over-smoothing. For example, while MPW is superior at capturing non-normal distributions, it produces kernels with too small of a bandwidth (hence the over-fitting), which is likely caused by the selected number of nearest neighbors used for the localized kernel estimation being too small. Meanwhile, compared to VC, ROME is numerically more stable and does not hallucinate new modes in the estimated densities (Figure 4). Furthermore, as part of several ablation studies, we found that ROME overcomes the shortcomings of other common density estimators, such as the over-fitting exhibited by kNN or the over-smoothing by GMM. In those studies, we additionally demonstrated that our approach of using clustering, decorrelation, and normalization is indispensable for overcoming the deficiencies of KDE.
@@ Acknowledgments
Contribution Statement
Anna Mészáros and Julian F. Schumann equally contributed to the design and implementation of the research, to the analysis of the results and to the writing of the manuscript and should be considered joint first authors. Javier Alonso-Mora provided valuable feedback on the writing of the manuscript. Arkady Zgonnikov and Jens Kober provided valuable feedback at all steps of the project and should be considered joint last authors.
@@ C Clustering Performance
$f_{\text{ROME}}$ $f_{\text{MPW}}$ $f_{\text{VC}}$
Aniso $\underline{0.007}\scriptscriptstyle{\pm 0.001}$ $0.016\scriptscriptstyle{\pm 0.001}$ $0.143\scriptscriptstyle{\pm 0.003}$
Varied $\underline{0.078}\scriptscriptstyle{\pm 0.003}$ $0.084\scriptscriptstyle{\pm 0.003}$ $0.087\scriptscriptstyle{\pm 0.003}$
@@
Normalization No normalization $f_{\text{GMM}}$
Decorrelation No decorrelation
Silhouette $0.007\scriptscriptstyle{\pm 0.001}$ $0.014\scriptscriptstyle{\pm 0.001}$ $0.039\scriptscriptstyle{\pm 0.001}$ $0.020\scriptscriptstyle{\pm 0.001}$ $0.054\scriptscriptstyle{\pm 0.001}$ $0.020\scriptscriptstyle{\pm 0.001}$ $0.001\scriptscriptstyle{\pm 0.000}$
DBCV $0.009\scriptscriptstyle{\pm 0.002}$ $0.014\scriptscriptstyle{\pm 0.001}$ $0.037\scriptscriptstyle{\pm 0.001}$ $0.022\scriptscriptstyle{\pm 0.001}$ $0.061\scriptscriptstyle{\pm 0.003}$ $0.023\scriptscriptstyle{\pm 0.002}$ $0.008\scriptscriptstyle{\pm 0.003}$
No clusters $0.095\scriptscriptstyle{\pm 0.001}$ $0.015\scriptscriptstyle{\pm 0.001}$ $0.245\scriptscriptstyle{\pm 0.002}$ $0.022\scriptscriptstyle{\pm 0.001}$ $0.039\scriptscriptstyle{\pm 0.001}$ $0.023\scriptscriptstyle{\pm 0.001}$ $0.530\scriptscriptstyle{\pm 0.002}$
@@
Silhouette $0.078\scriptscriptstyle{\pm 0.003}$ $0.083\scriptscriptstyle{\pm 0.003}$ $0.078\scriptscriptstyle{\pm 0.003}$ $0.083\scriptscriptstyle{\pm 0.003}$ $0.070\scriptscriptstyle{\pm 0.003}$ $0.083\scriptscriptstyle{\pm 0.003}$ $0.074\scriptscriptstyle{\pm 0.003}$
DBCV $0.074\scriptscriptstyle{\pm 0.003}$ $0.084\scriptscriptstyle{\pm 0.003}$ $0.074\scriptscriptstyle{\pm 0.003}$ $0.084\scriptscriptstyle{\pm 0.003}$ $0.065\scriptscriptstyle{\pm 0.003}$ $0.085\scriptscriptstyle{\pm 0.003}$ $0.068\scriptscriptstyle{\pm 0.003}$
No clusters $0.173\scriptscriptstyle{\pm 0.004}$ $0.081\scriptscriptstyle{\pm 0.003}$ $0.183\scriptscriptstyle{\pm 0.004}$ $0.080\scriptscriptstyle{\pm 0.003}$ $0.074\scriptscriptstyle{\pm 0.003}$ $0.080\scriptscriptstyle{\pm 0.003}$ $0.540\scriptscriptstyle{\pm 0.003}$
@@
Gaussian $\underline{0.006}\scriptscriptstyle{\pm 0.001}$ $0.023\scriptscriptstyle{\pm 0.002}$ $0.007\scriptscriptstyle{\pm 0.001}$ $-0.02\scriptscriptstyle{\pm 0.09}$ $-0.26\scriptscriptstyle{\pm 0.05}$ $\underline{-0.01}\scriptscriptstyle{\pm 0.11}$ $\underline{-2.84}\scriptscriptstyle{\pm 0.02}$ $-2.87\scriptscriptstyle{\pm 0.02}$ $-2.85\scriptscriptstyle{\pm 0.02}$
Elliptical $\underline{0.005}\scriptscriptstyle{\pm 0.001}$ $0.023\scriptscriptstyle{\pm 0.002}$ $0.007\scriptscriptstyle{\pm 0.001}$ $0.05\scriptscriptstyle{\pm 0.17}$ $-0.29\scriptscriptstyle{\pm 0.08}$ $\underline{0.04}\scriptscriptstyle{\pm 0.16}$ $\underline{-1.64}\scriptscriptstyle{\pm 0.02}$ $-1.66\scriptscriptstyle{\pm 0.02}$ $\underline{-1.64}\scriptscriptstyle{\pm 0.02}$
Rot. Ellip. $0.005\scriptscriptstyle{\pm 0.001}$ $0.023\scriptscriptstyle{\pm 0.002}$ $\underline{0.003}\scriptscriptstyle{\pm 0.001}$ $\underline{0.04}\scriptscriptstyle{\pm 0.20}$ $-0.35\scriptscriptstyle{\pm 0.11}$ $0.84\scriptscriptstyle{\pm 0.32}$ $\underline{-1.42}\scriptscriptstyle{\pm 0.02}$ $-1.45\scriptscriptstyle{\pm 0.02}$ $-1.55\scriptscriptstyle{\pm 0.01}$
Uni. Traj. $\underline{0.001}\scriptscriptstyle{\pm 0.000}$ $0.003\scriptscriptstyle{\pm 0.000}$ $0.801\scriptscriptstyle{\pm 0.003}$ $1.69\scriptscriptstyle{\pm 0.01}$ $2.33\scriptscriptstyle{\pm 0.01}$ $\underline{1.60}\scriptscriptstyle{\pm 0.11}$ $\underline{32.46}\scriptscriptstyle{\pm 0.02}$ $28.20\scriptscriptstyle{\pm 0.02}$ $-193.88\scriptscriptstyle{\pm 17.5}$
@@
Gaussian $\underline{0.004}\scriptscriptstyle{\pm 0.001}$ $0.012\scriptscriptstyle{\pm 0.001}$ $\underline{0.004}\scriptscriptstyle{\pm 0.001}$
Elliptical $\underline{0.004}\scriptscriptstyle{\pm 0.001}$ $0.012\scriptscriptstyle{\pm 0.001}$ $\underline{0.004}\scriptscriptstyle{\pm 0.001}$
Rot. Elliptical $\underline{0.004}\scriptscriptstyle{\pm 0.001}$ $0.013\scriptscriptstyle{\pm 0.001}$ $0.029\scriptscriptstyle{\pm 0.001}$
@@ E Comparison on Uni-modal Distribution
Norm. No norm. $f_{\text{GMM}}$
Clustering Decorr. No decorr.
Silhouette $0.004\scriptscriptstyle{\pm 0.001}$ $0.004\scriptscriptstyle{\pm 0.001}$ $0.017\scriptscriptstyle{\pm 0.001}$ $0.000\scriptscriptstyle{\pm 0.000}$
DBCV $0.004\scriptscriptstyle{\pm 0.001}$ $0.004\scriptscriptstyle{\pm 0.001}$ $0.017\scriptscriptstyle{\pm 0.001}$ $0.000\scriptscriptstyle{\pm 0.000}$
No clusters $0.004\scriptscriptstyle{\pm 0.001}$ $0.004\scriptscriptstyle{\pm 0.001}$ $0.017\scriptscriptstyle{\pm 0.001}$ $0.000\scriptscriptstyle{\pm 0.000}$
@@
Silhouette $0.004\scriptscriptstyle{\pm 0.001}$ $0.027\scriptscriptstyle{\pm 0.001}$ $0.038\scriptscriptstyle{\pm 0.001}$ $0.000\scriptscriptstyle{\pm 0.000}$
DBCV $0.004\scriptscriptstyle{\pm 0.001}$ $0.027\scriptscriptstyle{\pm 0.001}$ $0.038\scriptscriptstyle{\pm 0.001}$ $0.000\scriptscriptstyle{\pm 0.000}$
No clusters $0.004\scriptscriptstyle{\pm 0.001}$ $0.027\scriptscriptstyle{\pm 0.001}$ $0.038\scriptscriptstyle{\pm 0.001}$ $0.000\scriptscriptstyle{\pm 0.000}$
@@
Silhouette $\hphantom{0}32.46\scriptscriptstyle{\pm 0.02}$ $\hphantom{0}31.87\scriptscriptstyle{\pm 0.10}$ $-13.85\scriptscriptstyle{\pm 0.00}$ $\hphantom{0}28.61\scriptscriptstyle{\pm 0.02}$
DBCV $\hphantom{0}32.46\scriptscriptstyle{\pm 0.02}$ $\hphantom{0}31.87\scriptscriptstyle{\pm 0.10}$ $-14.51\scriptscriptstyle{\pm 1.96}$ $\hphantom{0}28.61\scriptscriptstyle{\pm 0.02}$
No clusters $\hphantom{0}32.46\scriptscriptstyle{\pm 0.02}$ $\hphantom{0}31.87\scriptscriptstyle{\pm 0.10}$ $-13.85\scriptscriptstyle{\pm 0.00}$ $\hphantom{0}28.61\scriptscriptstyle{\pm 0.02}$
@@
Silhouette $0.013\scriptscriptstyle{\pm 0.004}$ $0.166\scriptscriptstyle{\pm 0.063}$ $0.000\scriptscriptstyle{\pm 0.000}$ $0.005\scriptscriptstyle{\pm 0.003}$
DBCV $0.013\scriptscriptstyle{\pm 0.004}$ $0.166\scriptscriptstyle{\pm 0.063}$ $0.005\scriptscriptstyle{\pm 0.031}$ $0.005\scriptscriptstyle{\pm 0.003}$
No clusters $0.013\scriptscriptstyle{\pm 0.004}$ $0.166\scriptscriptstyle{\pm 0.063}$ $0.000\scriptscriptstyle{\pm 0.000}$ $0.005\scriptscriptstyle{\pm 0.003}$
@@
Silhouette $\hphantom{0}1.69\scriptscriptstyle{\pm 0.01}$ $\hphantom{0}1.68\scriptscriptstyle{\pm 0.02}$ $22.39\scriptscriptstyle{\pm 0.10}$ $\hphantom{0}2.20\scriptscriptstyle{\pm 0.03}$
DBCV $\hphantom{0}1.69\scriptscriptstyle{\pm 0.01}$ $\hphantom{0}1.68\scriptscriptstyle{\pm 0.02}$ $23.18\scriptscriptstyle{\pm 2.38}$ $\hphantom{0}2.20\scriptscriptstyle{\pm 0.03}$
No clusters $\hphantom{0}1.69\scriptscriptstyle{\pm 0.01}$ $\hphantom{0}1.68\scriptscriptstyle{\pm 0.02}$ $22.39\scriptscriptstyle{\pm 0.09}$ $\hphantom{0}2.20\scriptscriptstyle{\pm 0.03}$